4. The next three questions pertain to a region where there is a uniform magnetic field directed toward north.

On which of the following particles will the force exerted by the magnetic field be not zero?

a a proton moving toward south.

b an electron moving toward north.

c an electron moving toward east.

d a neutron moving toward south.

e a neutron moving toward west.
Answer #1

Correct option is : (c) An electron moving toward east.

Reason : Force exerted on a particle having charge q and velocity v, moving in a magnetic field of intensity B is F = qvB sin \theta , where, \theta is the angle between velocity vector ין and the magnetic field vector B .

So, to have a non - zero force, q must be non - zero, and, \theta must be not equal to 0o or 180o, so that, sin \theta remains non - zero.

For a proton or electron moving toward north, \theta = 0o, and, that moving toward south, \theta = 180o. So, F = 0.

Also, for a neutron, q = 0. So, F = 0.

Hence, the only option for having a non - zero magnetic force is an electron moving toward east, since, q is non - zero for an electron, and, \theta = angle between east and north = 90o in this case, so, sin \theta = sin 90o = 1.